### v2

# Auditing Student–AI Collaboration: A Case Study of Online Graduate CS Students

Abstract.
As generative AI becomes embedded in higher education, it increasingly shapes how students complete academic tasks. While these systems offer efficiency and support, concerns persist regarding over-automation, diminished student agency, and the potential for unreliable or hallucinated outputs. This study conducts a mixed-methods audit of student–AI collaboration preferences by examining the alignment between current AI capabilities and students’ desired levels of automation in academic work. Using two sequential and complementary surveys, we capture students’ perceived benefits, risks, and preferred boundaries when using AI. The first survey employs an existing task-based framework to assess preferences for and actual usage of AI across 12 academic tasks, alongside primary concerns and reasons for use. The second survey, informed by the first, explores how AI systems could be designed to address these concerns through open-ended questions. This study aims to identify gaps between existing AI affordances and students’ normative expectations of collaboration, informing the development of more effective and trustworthy AI systems for education.
AI in education, human–AI collaboration, student agency, automation preferences, generative AI, academic integrity, HCAI
copyright: acmlicensedjournalyear: 2026doi: 10.1145/XXXXXXX.XXXXXXXconference: Conference Full Name; June 03–05, 2026; Woodstock, NYbooktitle: Conference Full Name (Conference ’26), June 03–05, 2026, Woodstock, NYisbn: 978-1-4503-XXXX-X/26/06ccs: Human-centered computing Empirical studies in HCI
1. Introduction
Generative artificial intelligence (GenAI) tools such as ChatGPT, Claude, and Gemini are rapidly becoming embedded in higher education, reshaping how students complete academic tasks (Kasneci et al., 2023; Chan and Hu, 2024; Sukumar et al., 2025). These systems promise substantial gains in efficiency, personalization, and support, yet they also introduce significant risks related to over-automation, reduced student agency, and vulnerability to unreliable or hallucinated outputs (Shneiderman, 2020; Ji et al., 2023; Athaluri et al., 2023). As students increasingly delegate cognitive and procedural work to AI, it becomes essential to understand not only how they use these tools, but also how they perceive the trade-offs between automation and autonomy, and what design features would make AI systems more trustworthy in educational contexts (Khosravi et al., 2022; Amershi et al., 2019).
Recent research highlights the dual-edged nature of GenAI in education. Systematic reviews indicate that these tools can enhance writing quality, support personalized learning, and increase productivity, while simultaneously raising concerns about academic integrity, bias, and equitable access (Sukumar et al., 2025; Li et al., 2024; Camacho-Zúñiga et al., 2025). Empirical studies further suggest that the impact of AI depends heavily on how it is integrated pedagogically (Lee and Moore, 2024; Maizatul Akmar et al., 2025). However, a critical gap remains in understanding how students themselves navigate collaboration with AI across diverse academic tasks, particularly in graduate-level computer science education where technical accuracy and critical reasoning are paramount (Dwivedi et al., 2023; Sullivan et al., 2023).
This study addresses that gap by conducting a mixed-methods audit of student–AI collaboration among graduate students in the Georgia Tech Online Master of Science in Computer Science (OMSCS) program. Grounded in a Human-Centered AI (HCAI) perspective (Shneiderman, 2020, 2022) and informed by the Human Agency Scale (HAS) (Shao et al., 2025), we investigate two core dimensions of this collaboration. First, we measure students’ preferred levels of automation across twelve common academic tasks—spanning reading, writing, coding, studying, collaboration, and assessment—and compare these preferences to their actual AI usage (Shao et al., 2025, 2024b). This allows us to map tasks into four alignment zones: Green Light (high desire, high use), R&D Opportunity (high desire, low use), Low Priority (low desire, low use), and Red Light (low desire, high use). Second, through qualitative analysis of open-ended responses, we identify the system-level features—such as transparency, confidence indicators, explainability, hallucination warnings, and pedagogical alignment—that students believe would enhance the trustworthiness of AI in education (Amershi et al., 2019; Khosravi et al., 2022, 2024).
By integrating quantitative survey data with qualitative insights, this study provides a detailed, person-centered view of how graduate CS students perceive and negotiate the opportunities and risks of AI collaboration. Our findings contribute to the growing discourse on bidirectional human–AI alignment in education (Shen et al., 2024), offering evidence-based recommendations for educators, designers, and policymakers seeking to foster responsible, human-centered AI adoption in higher learning environments (Camacho-Zúñiga et al., 2025; Khosravi et al., 2024).
In this paper, we investigate the following research questions:
  • RQ1: What levels of automation do students desire across different academic tasks, and how do these preferences align with their actual AI usage?
  • RQ2: How do students’ reasons for using AI and their associated concerns vary across task types, and what does this reveal about the perceived risks and benefits of student–AI collaboration?
  • RQ3: What system features and design principles do students identify as most important for addressing their concerns ?
To answer these questions, we administered a two-part survey to over 50 OMSCS students. Part 1 employed a task-based adaptation of the Human Agency Scale (Shao et al., 2025) to quantify automation preferences and actual usage, as well as students’ primary reasons for and concerns about AI use across tasks. Part 2 used an open-ended prompt to elicit qualitative feedback on features that would address these concerns in educational AI (Braun and Clarke, 2006). Our analysis integrates these data to produce an automation alignment map and a thematic framework for trustworthy AI design.
Our findings indicate that students strongly desire automation for repetitive, time-consuming, or accuracy-focused tasks, but remain cautious about delegating activities that involve higher-order reasoning, creativity, or technical nuance (Cotton et al., 2023; Halaweh, 2023). Notably, no tasks fell into the Red Light Zone, suggesting that current usage does not exceed students’ comfort levels. Students also consistently called for greater transparency, confidence indicators, source citations, and safeguards against hallucinations—preferences that echo prior calls for explainable and controllable AI systems (Amershi et al., 2019; Khosravi et al., 2022, 2024).
This research makes three main contributions:
  • Empirical: We provide a detailed, mixed-methods audit of student–AI collaboration in a real-world graduate CS context, capturing both behavioral patterns and perceived trust factors.
  • Methodological: We demonstrate the utility of the Human Agency Scale (Shao et al., 2025) and alignment zone mapping for evaluating automation acceptance in education.
  • Practical: We offer evidence-based design implications and policy recommendations for creating more transparent, reliable, and human-centered AI tools in higher education (Amershi et al., 2019; Shneiderman, 2022).
2. Related Work
Research on artificial intelligence in education has expanded rapidly with the emergence of large-scale generative language models. Prior work spans educational research, learning sciences, human–computer interaction, and socio-technical studies of automation, addressing both opportunities and risks of AI-supported learning. This section synthesizes three strands most relevant to our study: (1) empirical and ethical research on generative AI in higher education, (2) Human-Centered AI (HCAI) as a design framework for trustworthy educational technologies, and (3) approaches to measuring human agency and automation preferences, emphasizing the Human Agency Scale (HAS) as a recent operationalization for quantifying desired human involvement.
2.1. Generative AI in Higher Education
A growing body of literature documents the transformative but contested role of generative AI in academic settings. Systematic reviews and meta-analyses highlight that tools such as ChatGPT can improve writing fluency, reduce cognitive load for routine tasks, and support personalized learning experiences when used appropriately (Kasneci et al., 2023). These benefits are most apparent for surface-level tasks like grammar correction, summarization, and initial drafting, where AI functions as a productivity aid rather than a substitute for learning.
At the same time, substantial concerns have been raised about over-reliance on AI, hallucinated or biased outputs, and the potential erosion of critical thinking and disciplinary epistemic skills (Bastani et al., 2024; Selwyn, 2019). Empirical work suggests that when students delegate substantive intellectual work—such as complex problem solving or argumentation—to AI systems without appropriate scaffolding, learning outcomes can suffer (Bastani et al., 2024). These risks are particularly salient in technically demanding disciplines such as computer science, where correctness, abstraction, and conceptual understanding are central to education.
Instructional context and pedagogical design play a decisive role in mediating these effects. For example, research comparing AI-generated feedback with human or peer feedback finds that AI is effective for addressing surface-level concerns aligned with rubrics, but that peer or instructor feedback remains essential for fostering deeper conceptual understanding and disciplinary reasoning (Cho and Schunn, 2007; Liu et al., 2024). Such work converges on the conclusion that generative AI is most beneficial when integrated as a complement to human pedagogical interaction, rather than as a replacement.
In addition to learning outcomes, generative AI raises fundamental challenges for academic integrity, authorship, and equity. Scholars argue that traditional assessment frameworks are increasingly misaligned with AI-enabled workflows, driving the need for new assessment designs and integrity policies (Perkins, 2023b). Disparities in access to AI tools and differences in AI literacy further risk exacerbating educational inequalities, reinforcing calls for responsible, inclusive integration of AI in education (Holmes et al., 2022). Together, this literature underscores the need for approaches that harness efficiency gains while preserving autonomy, integrity, and equitable participation.
2.2. Human-Centered AI in Educational Contexts
Human-Centered AI (HCAI) has emerged as a foundational paradigm for responding to the risks posed by opaque, overly autonomous AI systems. Rooted in human–computer interaction and AI ethics, HCAI advocates for systems that are reliable, safe, and trustworthy while preserving meaningful human control and agency (Shneiderman, 2020). Rather than optimizing for automation alone, HCAI emphasizes augmentation that supports human goals, values, and accountability.
In educational contexts, HCAI principles foreground the importance of preserving both student and instructor agency, enhancing transparency of system behavior, and embedding safeguards against misuse, bias, and automation bias (Amershi et al., 2019; Holmes et al., 2022). Design recommendations from this literature include making system limitations visible, offering explanations and confidence cues, and enabling users to intervene, override, or verify AI outputs. These features are particularly critical in learning environments where inappropriate trust or over-dependence on AI can undermine motivation, metacognition, and skill development.
A concept in HCAI is bidirectional alignment: AI systems should be designed to reflect human values and pedagogical aims, while users must develop the critical capacity to engage with AI systems responsibly (Shen et al., 2024). In educational settings, this mutual adaptation matters because student–AI interactions can shape epistemic beliefs, study approaches, and perceptions of authorship and ownership of work. Our work adopts this perspective by conceptualizing student–AI collaboration as an ongoing partnership requiring thoughtful system design and active human engagement.
2.3. Human Agency and Automation Preferences
As AI systems become active collaborators in human workflows, researchers have sought to quantify how users perceive and negotiate their agency under automation. A recent contribution to this literature is the **Human Agency Scale (HAS)** introduced by Shao et al., which provides a five-level scale (H1–H5) to quantify the degree of human involvement preferred in task performance when working with AI agents (Shao et al., 2025). HAS moves beyond binary automate-or-not frameworks by offering a shared language for examining desired human control versus automation or augmentation across diverse tasks and occupations.
The HAS differentiates between complete human control (higher levels) and higher degrees of automation (lower levels), allowing for nuanced analyses of user preferences in human–AI work contexts. Validation and application of HAS indicate that agency perceptions vary systematically with task criticality, trust, perceived AI competence, and potential consequences of error, making the scale well-suited to assess collaborative human–AI workflows. Related research on algorithm aversion and resistance to automation shows that users are often reluctant to delegate control in domains tied to expertise, identity, or accountability (Dietvorst et al., 2015; Longoni et al., 2019).
Within educational research, HAS-aligned findings suggest that students are typically comfortable delegating repetitive or mechanical tasks — such as formatting, grammar correction, or simple content summarization — while preferring to retain control over tasks that involve creativity, reasoning, or deep disciplinary judgment (Liu et al., 2024; Bastani et al., 2024). These preferences are shaped by trust, perceived risk, and confidence in one’s ability to evaluate AI outputs — critical factors that influence appropriate reliance on automated systems.
Our study adapts the HAS framework to graduate computer science education, applying it at a task-specific level to compare desired human agency, actual AI use, and confidence in verifying outputs. This approach enables a fine-grained audit of student–AI collaboration and provides empirical grounding for human-centered design recommendations.
2.4. Research Gap
Taken together, prior work establishes the promise of generative AI for education, the pedagogical and ethical risks it poses, and the importance of agency-preserving, human-centered design. However, much existing research focuses either on macro-level outcomes (e.g., policy debates, institutional assessment practices) or on system-centric evaluations of AI performance. There is limited attention to how students themselves experience and negotiate collaboration with AI across concrete academic tasks, especially in disciplines with strong technical demands.
Few studies integrate quantitative measures of automation preference with qualitative insights into trust, transparency, and desired system features. By conducting a mixed-methods audit of student–AI collaboration grounded in HCAI principles and operationalized through the Human Agency Scale, our work addresses this gap and contributes a person-centered perspective to the design, adoption, and governance of AI in higher education.
3. Methodology
Table 1. 12 academic tasks used in the study.
ID Task Description
T1 Summarize a research article.
T2 Formatting Citation and Bibliography.
T3 Brainstorm essay ideas or outline structure
T4 Revise Writing for grammar and style.
T5 Debug code snippet or explain errors.
T6 Create personalized study plan.
T7 Generate flashcards or practice quizzes
T8 Recommend Learning Resources such as video or textbook
T9 Draft professional emails to TA or Professor
T10 Help explain or solve step by step quantitative problem
T11 Summarize class note or discussion
T12 Give feedback on draft writing
3.1. 12 Academic Tasks
To systematically evaluate student preferences for AI collaboration, we selected twelve representative academic tasks that span the core cognitive and procedural activities common in graduate-level computer science education (Swiecki et al., 2022). These tasks were drawn from six functional categories: (1) Reading and Note-Making, (2) Writing and Revision, (3) Coding and Problem Solving, (4) Studying and Metacognition, (5) Collaboration and Communication, and (6) Assessment and Feedback. This categorization reflects the multifaceted nature of graduate learning, where students engage not only in content acquisition but also in synthesis, creation, and critical evaluation (Kasneci et al., 2023).
Each task was chosen based on its relevance to both conceptual and practical skill development. For instance, tasks such as debugging code (T5) and solving quantitative problems step-by-step (T10) represent core technical competencies, while activities like summarizing research (T1) and providing feedback on writing (T12) emphasize critical engagement and communication skills (Shao et al., 2024a). By including tasks that range from routine (e.g., formatting citations) to complex and open-ended (e.g., brainstorming essay structures), we aimed to capture a spectrum of automation preferences and perceived risks.
The tasks were designed to vary along several dimensions relevant to AI collaboration:
Cognitive Demand: From low-level, rule-based tasks (T2, T4) to high-level, interpretive tasks (T3, T12) (Swiecki et al., 2022).
Structured vs. Open-Ended: Well-defined tasks with clear correctness criteria (T5, T10) versus tasks requiring creativity or subjective judgment (T3, T6) (Mollick and Mollick, 2023).
Frequency and Time Cost: Common, time-consuming activities (T1, T11) versus occasional but high-stakes tasks (T9, T12).
Potential for AI Error: Tasks where AI hallucinations or inaccuracies could have meaningful consequences (T1, T5, T10) (Ji et al., 2023; Athaluri et al., 2023).
This variation allows us to examine how task characteristics shape students’ willingness to delegate work to AI, their confidence in verifying outputs, and their primary concerns regarding reliability and learning integrity (Cotton et al., 2023; Halaweh, 2023).
Table 1 presents the complete list of tasks.
3.2. Desired Automation Likert Questions
Likert Question for Collecting Automation Desire Question. For each academic task, please rate the level of automation that an AI system should ideally provide to support learning, from your (a student’s) perspective. Scale. 1: AI should not get involved in this task at all 2: AI may give minor suggestions 3: AI and human should share responsibility 4: AI should handle most of the task with human oversight 5: AI should perform this task autonomously
Figure 1. Illustration of the Likert-scale question used to measure students’ desired level of AI automation for academic tasks.
Following prior work by Shao et al. (Shao et al., 2025) on characterizing levels of AI automation, we operationalize students’ automation preferences using a five-point Likert-scale that reflects a continuum of responsibility allocation between humans and AI (Figure 1). Rather than treating automation as a binary choice, this scale explicitly distinguishes between AI serving a purely assistive role, sharing responsibility with the human learner, or assuming primary or full control over the task (Shao et al., 2024b). This design allows us to capture nuanced automation boundaries across academic tasks and to examine where students draw limits on AI involvement in educational contexts (Venkatesh et al., 2003).
3.3. Actual Usage Likert Question
Likert Question for Observed AI Usage Question. For each academic task listed below, please indicate how capable current AI systems are based on your own experience, reflected by how you actually use AI for that task. Scale. 1: I do not use AI for this task at all 2: I use AI to provide minor or partial help 3: I use AI for this task moderately 4: AI handles most of the task with human oversight 5: I rely entirely on AI to automate this task
Figure 2. Illustration of the Likert-scale question used to capture students’ self-reported observed AI usage across academic tasks.
In contrast to prior work by Shao et al. (Shao et al., 2025), which characterizes AI capability using expert annotations, our study does not rely on external expert judgments of system performance. Instead, we capture students’ perceived AI capability through their self-reported actual usage and reliance on AI systems across academic tasks (Figure 2(Dwivedi et al., 2023; Sullivan et al., 2023).
Specifically, we ask participants to indicate how they currently use AI for each task, ranging from not using AI at all to relying entirely on AI to automate the task. This formulation operationalizes perceived capability in terms of habitual use and practical reliance, reflecting students’ lived experiences with contemporary AI tools rather than an objective assessment of model correctness or task success (Jian et al., 2000).
While self-reported usage does not substitute for expert evaluation, it provides a meaningful proxy for how capable AI systems are perceived to be in real educational settings, particularly in contexts where students make day-to-day decisions about whether and how much to rely on AI. Maintaining the same automation continuum as the desired automation measure allows us to directly compare normative expectations (how AI should be used) with descriptive practices (how AI is actually used).
3.4. Four-Zone Classification of AI Automation
To analyze students’ perceptions of appropriate AI involvement in academic tasks, we adopted a four-zone classification of automation adapted from Shao et al.’s framework (Shao et al., 2025) for integrating human worker and AI expert perspectives on automation. This framework partitions the automation landscape into four zones based on the alignment or misalignment between human preferences and perceived AI capability.
The four zones are defined as follows:
Automation “Green Light” Zone. Tasks in the Green Light zone are those for which both humans and AI systems are perceived as well-suited for automation. In this zone, there is strong alignment between participants’ willingness to delegate a task to AI and their confidence in AI’s ability to perform the task effectively. These tasks represent contexts in which AI adoption is broadly acceptable and likely to provide immediate benefits with minimal resistance (Shao et al., 2025).
Automation “Red Light” Zone. The Red Light zone includes tasks that participants believe should not be automated, even if AI systems are technically capable of performing them. This zone reflects strong human resistance to AI involvement, often due to concerns related to trust, accountability, ethical implications, or academic integrity (Cotton et al., 2023; Perkins, 2023a). Tasks in this zone signal boundaries where AI use is perceived as inappropriate despite potential performance gains.
R&D Opportunity Zone. Tasks in the R&D Opportunity zone are those for which participants express openness to AI assistance but lack confidence in current AI systems’ capabilities. This zone highlights gaps between user expectations and existing technological performance and points to opportunities for further AI development, evaluation, and design refinement (Shao et al., 2025). Improving transparency, reliability, or controllability of AI systems may enable tasks in this zone to transition toward the Green Light zone (Kasneci et al., 2023).
Low Priority Zone. The Low Priority zone consists of tasks that participants neither desire to automate nor believe AI systems are well-suited to perform. These tasks are typically viewed as either unimportant for automation or inherently human-centered, resulting in low perceived value of AI involvement. From a design perspective, this zone suggests limited benefit in prioritizing AI development for such tasks (Shao et al., 2025).
In our study, participants’ Likert-scale responses regarding desired levels of AI automation and confidence in AI-generated outputs were jointly considered to map academic tasks onto these four zones. This classification enabled a systematic comparison of where students perceive AI use as appropriate, inappropriate, promising but underdeveloped, or low priority within the context of AI-assisted coursework.
3.5. Primary Concerns and Reasons for AI Usage across Academic Tasks
Category Survey Item Description
Primary Concerns
Inaccurate or misleading information Concern that AI outputs may contain factual errors or hallucinations that affect task correctness
Risk of academic misconduct Concern about plagiarism, policy violations, or inappropriate AI use in graded work
Reduced critical thinking Concern that relying on AI may weaken students’ independent reasoning or learning
Reasons for AI Usage
Saving time Using AI to complete tasks more efficiently and reduce time spent on repetitive work
Reducing cognitive load Using AI to manage mentally demanding or complex tasks
Improving output quality Using AI to enhance clarity, accuracy, or polish of academic outputs
Table 2. Selection of primary concerns and reasons for AI usage included in the survey.
To understand what shapes students’ adoption of AI across different academic tasks, we asked participants to report both their primary concerns when using AI and their reasons for relying on AI assistance (Dwivedi et al., 2023; Sullivan et al., 2023). Concerns included risks such as inaccurate or misleading outputs, hallucinations, academic misconduct, and reduced critical thinking (Cotton et al., 2023; Halaweh, 2023; Imran and Almusharraf, 2023), while reasons for use captured perceived benefits such as saving time, reducing cognitive load while managing conceptually complex task and improving accuracy or polishing academic outputs (Kasneci et al., 2023).
Analyzing these responses at the task level allows us to examine how motivations for AI use coexist with, and are constrained by, students’ concerns. While students often report using AI to support efficiency and reduce effort in repetitive or demanding tasks, concerns about hallucinations and reliability remain salient, particularly for tasks involving complex reasoning or high academic stakes (Ji et al., 2023; Athaluri et al., 2023; Weiser, 2023). This dual perspective highlights that students’ AI use is not driven solely by convenience, but reflects ongoing trade-offs between perceived benefits and epistemic risks (Bender et al., 2021).
These findings provide important context for our subsequent analysis, motivating a closer examination of how students experience AI hallucinations in practice and what system-level design features they expect to better address these concerns.
3.6. Design Expectations for Addressing AI Concerns
In addition to structured Likert-scale items, we included open-ended questions to elicit students’ perspectives on how AI systems should be designed to address their concerns in educational contexts (Braun and Clarke, 2006). These questions focused on design features and interaction mechanisms that could improve trustworthiness, support error detection, and mitigate risks such as hallucination, without constraining responses to predefined options (Perkins, 2023a).
The use of open-ended questions allows participants to express design expectations in their own terms, capturing aspects of AI system behavior and interface design that may not be fully anticipated by closed-form survey items (Braun and Clarke, 2006). We treat these responses as qualitative inputs for identifying recurring design considerations and informing subsequent analysis of user-centered approaches to AI alignment in education (Swiecki et al., 2022).
4. User Study Method
4.1. Participants
Participants were graduate students enrolled in the Online Master of Science in Computer Science (OMSCS) program at the Georgia Institute of Technology, specifically students taking CS6460: Educational Technology. Participation in the survey was voluntary and counted toward a course participation grade.
To ensure the relevance of responses to AI-assisted coursework, we filtered out participants who reported rarely using or not using AI tools in their academic work. In total, $N=57$ students completed the survey. After filtering, $N=44$ participants who reported at least occasional use of AI systems for coursework were retained for analysis. The final sample therefore consists of students with meaningful experience interacting with AI in educational contexts, allowing us to focus our analysis on substantive patterns of AI use.
4.2. Survey Based Interviews
To examine students’ perceptions and experiences with AI-assisted coursework, we employed a survey that combined Likert-scale questions with open-ended response items. Rather than conducting separate interviews, the questionnaire itself served as the primary data collection instrument, allowing participants to provide both structured ratings and written explanations of their views.
The Likert-scale questions asked participants to evaluate their use of AI tools (e.g., frequency of use), their confidence in assessing AI-generated outputs, and their perceptions of the appropriate level of AI automation across a range of academic tasks, including writing, programming, problem solving, and studying. These items enabled quantitative analysis of general trends in students’ attitudes toward AI use in coursework.
Complementing the Likert-scale items, the survey included open-ended questions that prompted participants to elaborate on their responses. These questions asked participants to explain their reasoning behind their ratings, describe how they currently use AI tools for specific academic tasks, and articulate any concerns they have regarding accuracy, bias, over-reliance, or academic integrity when using AI systems. The open-ended responses thus functioned as survey-based interviews, providing qualitative insight into participants’ perspectives in their own words.
All responses were collected anonymously.The qualitative data from the open-ended questions were analyzed using an iterative thematicanalysis approach, beginning with open coding to identify recurring concepts,followed by axial coding to refine and organize emergent themes. This mixed-format questionnaire design allowed us to triangulate quantitative trends with qualitative explanations, strengthening the interpretability of our findings.
4.3. Ethics
The study followed established ethical guidelines for research involving human participants. Prior to participation, all participants completed an informed consent form. Participation was voluntary, and no personally identifiable information was collected. All responses were anonymized prior to analysis and used solely for research purposes.
The study posed minimal risk to participants. All questions focused on participants’ experiences and perceptions of AI use in academic contexts, and no deception was involved. Data were collected and stored in a manner that protected participant privacy and ensured confidentiality.
ID Age Group OMSCS Specialization AI Usage Frequency
P1 30–40 Human–Computer Interaction several times per week
P2 25–30 Machine Learning Daily or almost daily
P3 25–30 Human–Computer Interaction Often
P4 18–24 Artificial Intelligence Often
P5 30–40 Human–Computer Interaction several times per week
P6 30–40 Artificial Intelligence several times per week
P7 above 40 Computing Systems several times per week
P8 30–40 Artificial Intelligence Daily or almost daily
P9 18–24 Artificial Intelligence several times per week
P10 18–24 Not specified Daily or almost daily
P11 18–24 Artificial Intelligence Daily or almost daily
P12 18–24 Not selected yet Often
P13 30–40 Not decided yet Daily or almost daily
P14 30–40 Human–Computer Interaction Daily or almost daily
P15 18–24 Undecided several times per week
P16 30–40 Computing Systems Often
P17 25–30 Artificial Intelligence Often
P18 30–40 Human–Computer Interaction several times per week
P19 25–30 Computing Systems several times per week
P20 18–24 Human–Computer Interaction several times per week
P21 30–40 Artificial Intelligence Daily or almost daily
P22 18–24 Artificial Intelligence Often
P23 30–40 Machine Learning Daily or almost daily
P24 25–30 Human–Computer Interaction Daily or almost daily
P25 25–30 Computing System Daily or almost daily
P26 30–40 Artificial Intelligence Daily or almost daily
P27 18–24 Machine Learning Daily or almost daily
P28 18–24 Artificial Intelligence / Machine Learning several times per week
P29 above 40 Artificial Intelligence Daily or almost daily
P30 18–24 Undecided several times per week
P31 25–30 Artificial Intelligence Daily or almost daily
P32 18–24 Artificial Intelligence Daily or almost daily
P33 18–24 Not specified Daily or almost daily
P34 25–30 Human–Computer Interaction several times per week
P35 18–24 Artificial Intelligence Often
P36 18–24 Artificial Intelligence Often
P37 30–40 Computing System Daily or almost daily
P38 25–30 Human–Computer Interaction Daily or almost daily
P39 25–30 Human–Computer Interaction several times per week
P40 above 40 Machine Learning Daily or almost daily
P41 18–24 Artificial Intelligence several times per week
P42 25–30 Human–Computer Interaction several times per week
P43 25–30 Human–Computer Interaction Daily or almost daily
P44 30–40 Artificial Intelligence Daily or almost daily
Table 3. Participant demographics of OMSCS students who reported at least occasional use of AI tools in coursework.
4.4. Data Analysis
To compare students’ desired levels of AI involvement with their reported use of AI tools, we operationalized two aggregate measures: aggregate desire level and aggregate usage level. Both measures were derived from participants’ Likert-scale responses and computed at the task level.
Aggregate Desire Level.
For each academic task $t$, participants indicated the level of AI automation they believed was appropriate using an ordered Likert-scale corresponding to increasing levels of automation (from minimal AI involvement to fully autonomous AI execution). Responses were numerically encoded such that higher values indicate a greater desired level of AI automation. The aggregate desire level for task $t$ was calculated as the mean of these encoded responses across all participants in the analytic sample ($N=44$):
$D_{t}=\frac{1}{N}\sum_{i=1}^{N}d_{i,t},$
where $d_{i,t}$ denotes participant $i$’s desired automation rating for task $t$.
Aggregate Usage Level.
Actual AI usage was operationalized using participants’ self-reported frequency of AI tool use. Responses were encoded on an ordered numerical scale, with higher values indicating more frequent AI use. The aggregate usage level was computed as the mean usage score across all participants:
$U=\frac{1}{N}\sum_{i=1}^{N}u_{i},$
where $u_{i}$ denotes participant $i$’s reported frequency of AI use.
When task-specific usage data were available, aggregate usage was computed analogously at the task level:
$U_{t}=\frac{1}{N}\sum_{i=1}^{N}u_{i,t}.$
5. User Study Result
Figure 3. Comparison of students’ desired AI automation levels and self-reported actual AI usage across twelve academic tasks (T1–T12).
Figure 4. Four-zone automation alignment map plotting desired AI automation against actual AI usage for twelve academic tasks.
5.1. RQ1: Desired Automation Level vs. Actual AI Usage Level
Figure 3 illustrates participants’ aggregate desired automation levels and reported AI usage levels across twelve academic tasks (T1–T12). Across most tasks, desired automation levels exceed reported usage, indicating a consistent gap between participants’ preferences for AI involvement and their current patterns of AI use. However, the magnitude of this gap varies substantially by task type, suggesting that task characteristics play an important role in shaping how students engage with AI tools in practice.
5.1.1. Information Compression and Retrieval Tasks
Tasks involving summarization and information organization (T1: summarizing research articles; T11: summarizing class notes or discussions) exhibit relatively high desired automation levels alongside moderately high actual usage. These tasks are primarily extractive and supportive in nature, requiring limited original reasoning from the student. Participants appear relatively comfortable relying on AI for such tasks, likely because errors are easier to identify and correct, and inaccuracies are perceived as less consequential. Accordingly, the gap between desired automation and actual usage is comparatively small for these tasks.
5.1.2. Writing Support and Surface-Level Editing Tasks
Tasks such as revising writing for grammar and style (T4), formatting citations and bibliographies (T2), and giving feedback on draft writing (T12) show some of the highest desired automation levels across all tasks. These activities are procedural and rule-based, making them well suited for AI assistance in principle. Nevertheless, actual usage remains consistently lower than desired. This gap suggests that, despite recognizing AI’s potential utility, participants may exercise caution in practice, possibly due to concerns about correctness, formatting conventions, or adherence to academic norms.
5.1.3. Ideation and Study Support Tasks
Tasks related to brainstorming essay ideas (T3), creating personalized study plans (T6), generating flashcards or practice quizzes (T7), and recommending learning resources (T8) display some of the largest gaps between desired automation and actual usage. While participants express strong interest in AI support for these tasks, their reported usage remains comparatively low. These tasks require contextual understanding, personalization, and alignment with individual learning goals, which participants may perceive as areas where current AI tools do not yet consistently meet their expectations. The pronounced gap reflects openness to AI assistance paired with reservations about its present capabilities.
5.1.4. Communication and Socially Sensitive Tasks
Drafting professional emails to instructors or teaching assistants (T9) shows a moderate desired automation level but relatively lower reported usage compared to many other tasks. Unlike more technical activities, this gap appears to be driven less by perceptions of AI capability and more by social and interpersonal considerations. Participants may be cautious about delegating tasks that involve tone, professionalism, and accountability, preferring to retain greater human control even when AI assistance could be beneficial.
5.1.5. Analytical and High-Stakes Reasoning Tasks
Analytical tasks exhibit heterogeneous patterns. For debugging code (T5), participants report relatively high AI usage, with usage levels closely tracking desired automation. This suggests that students are already comfortable using AI tools to support code-related tasks, likely because outputs can be readily validated through testing and execution. In contrast, step-by-step quantitative problem solving (T10) shows a larger gap between desired automation and actual usage. Although participants express moderate interest in AI support for such tasks, their reported usage remains lower. These tasks demand precise reasoning and conceptual understanding, and errors may be more difficult to detect, leading participants to exercise greater caution in relying on AI assistance.
5.1.6. Four Automation Zones
Figure 4 shows that the twelve academic tasks are distributed unevenly across the four automation zones, revealing systematic patterns in how students’ preferences for AI automation align with their reported usage.
Several tasks cluster in the region characterized by both relatively high desired automation and high actual AI usage. These include formatting citations and bibliographies, revising writing for grammar and style, summarizing class notes or discussions, summarizing research articles, and debugging code. The concentration of tasks in this region suggests that students are already actively using AI for tasks that are procedural, supportive, or whose outputs can be readily verified.
A second group of tasks falls into a region where desired automation is high but actual usage remains comparatively lower. This group includes creating personalized study plans, generating flashcards or practice quizzes, recommending learning resources, drafting professional emails, and solving step-by-step quantitative problems. The placement of these tasks indicates that while students express interest in greater AI involvement, their current usage lags behind their preferences, highlighting areas where existing AI tools may not yet fully meet students’ expectations or where contextual concerns limit adoption.
Only a small number of tasks appear in regions associated with both low desired automation and low actual usage. Brainstorming essay ideas or outlining structure is the most prominent example, suggesting that participants prefer to retain human control over tasks they view as central to creativity or personal thinking.
5.2. RQ2: Primary Reasons and Concerns of using AI across academic tasks
Figure 5. Distribution of students’ reported reasons for using AI across twelve academic tasks.
Figure 6. Distribution of students’ reported concerns regarding AI use across twelve academic tasks. Percentages do not sum to 100% because the survey included an explicit “All of them” option, which could be selected by participants.
Figures 5 illustrate the distributions of students’ reported reasons for using AI, and their associated concerns across twelve academic tasks was illustrated in Figures 6.
5.2.1. Writing and Revision Tasks
Writing- and revision-oriented tasks (T1: summarizing research articles; T4: revising writing for grammar and style; T9: drafting professional emails; T11: summarizing class notes or discussions; T12: giving feedback on draft writing) exhibit reason distributions that are strongly dominated by efficiency-related motivations. Across most tasks in this category, saving time accounts for the largest proportion of selected reasons, typically exceeding half of all responses. This pattern reflects students’ perception of writing-related AI use as a form of pragmatic assistance that reduces routine effort rather than fundamentally reshaping the intellectual substance of the task. In addition, improving output quality constitutes a substantial secondary share, particularly for revision-focused tasks such as T4 and T12, where AI is viewed as effective in enhancing clarity, grammatical correctness, and surface-level polish.
The combination of these motivations suggests that students conceptualize AI as a tool that complements existing writing practices rather than replaces them. Tasks in this category generally involve outputs that are readily inspectable, editable, and attributable to the student, which may further lower perceived barriers to AI adoption. As a result, students appear comfortable leveraging AI to accelerate drafting or refinement while retaining ultimate control over content and intent.
Concern distributions for writing and revision tasks are comparatively benign. For several activities (e.g., T9 and T11), None represents the largest proportion of reported concerns, indicating that many students perceive little risk in using AI for these purposes. When concerns do arise, they are most frequently associated with inaccurate information, which consistently outweighs concerns related to reduced critical thinking or plagiarism. This suggests that students’ primary vigilance lies in factual correctness rather than in broader ethical or cognitive implications. Taken together, these patterns indicate a relatively high level of trust in AI-assisted writing support, conditioned on students’ ability to verify and revise outputs as needed.
5.2.2. Ideation and Planning Tasks
Ideation and planning tasks (T3: brainstorming essay ideas or outlining structure; T6: creating a personalized study plan) demonstrate more differentiated and nuanced distributions of both reasons and concerns. For brainstorming (T3), motivations are more evenly distributed across categories, with improving output quality and reducing cognitive load together accounting for a substantial proportion of responses. This reflects students’ interest in AI as a means of generating ideas, exploring alternative structures, and overcoming initial barriers in the early stages of composition.
However, the corresponding concern distribution for T3 reveals a sharply contrasting pattern. Reduced critical thinking overwhelmingly dominates the concern profile, far exceeding all other categories. This divergence highlights a strong awareness among students of the cognitive risks associated with delegating ideation and creative reasoning to AI. While AI is valued for sparking ideas or providing structure, students appear wary of over-reliance in tasks that are closely tied to learning outcomes, originality, and intellectual ownership. The coexistence of strong perceived benefits and pronounced concerns suggests that ideation tasks occupy a particularly sensitive boundary between assistance and substitution.
In contrast, study planning (T6) exhibits a more straightforward profile. Here, saving time emerges as the dominant reason, while None constitutes the largest share of concerns. This indicates that students perceive AI-supported planning as primarily organizational rather than cognitively generative. Because study planning focuses on scheduling, resource allocation, and logistical coordination, AI assistance in this context is seen as unlikely to undermine deep understanding or critical engagement. The contrast between T3 and T6 underscores students’ ability to differentiate between tasks that involve creative reasoning and those that primarily involve coordination and efficiency.
5.2.3. Learning Support Tasks
Learning support tasks (T7: generating flashcards or practice quizzes; T8: recommending learning resources) are characterized by highly concentrated reason distributions. In both cases, saving time constitutes the clear majority of responses, with improving output quality playing a secondary role. These distributions indicate that students primarily view AI as a means of accelerating access to study materials and reducing the overhead associated with preparation and review.
Concern distributions for these tasks further reinforce this interpretation. Across both T7 and T8, None consistently represents the largest share of concerns, while accuracy- and cognition-related concerns remain comparatively minor. This suggests that students perceive these tasks as low stakes, with AI functioning as a supplementary aid rather than a substitute for core learning or evaluative processes. Because outputs such as flashcards or resource recommendations can be selectively used, ignored, or verified, students appear comfortable integrating AI with minimal hesitation. Overall, these patterns position AI as an efficiency-enhancing study aid that supports learning without threatening autonomy or understanding.
5.2.4. Technical and Quantitative Problem-Solving Tasks
Technical and analytical tasks (T5: debugging code or explaining programming errors; T10: step-by-step quantitative problem solving) exhibit the most balanced and complex distributions of reasons and concerns. For both tasks, reducing cognitive load accounts for the largest proportion of reported reasons, highlighting students’ reliance on AI to manage complexity, interpret errors, and scaffold multi-step reasoning. Improving output quality also represents a meaningful share, reflecting the importance of correctness and precision in technical contexts.
At the same time, concern distributions reveal elevated perceived risks. For both T5 and T10, inaccurate information and reduced critical thinking together constitute a substantial proportion of reported concerns, exceeding those observed in other task categories. This pattern suggests that while students value AI assistance for navigating difficult technical material, they remain acutely aware of the potential consequences of errors and the risk of superficial understanding. Unlike writing or learning support tasks, mistakes in technical or quantitative domains may be harder to detect and more costly in terms of learning outcomes. The coexistence of strong efficiency motivations and prominent concerns therefore reflects a cautious, evaluative stance toward AI use, where benefits are actively weighed against the risks of dependency and misunderstanding in high-stakes analytical contexts.
5.3. RQ3: Expected System Features and Design Principles of AI System
After participants reported their primary reasons and concerns regarding AI use across academic tasks, we administered a follow-up survey to collect open-ended responses. This resulted in 53 valid responses ($N=53$), in which participants described the features and design principles they expect AI systems to provide to address concerns about inaccurate information and hallucinations. We refer to individual respondents as R1–R53 throughout this section.
Table 4. Students’ expected features and design principles for educational AI systems.
Reason Example quotes Potential benefits
Transparency and verifiability of AI outputs “Always returning a clickable link to a source so it can be easily validated.” (R10); “Giving sources for all the information they give.” (R22); “AI systems should always provide where it’s getting its data from.” (R42) Students can independently verify AI-generated information and assess its reliability, reducing blind trust and inappropriate academic use.
Explicit communication of uncertainty “It should have a percentage next to every response.” (R23); “A certainty factor would be an interesting addition.” (R4); “More transparency about uncertainty or when guessing is used.” (R34) Students can better calibrate trust in AI outputs and decide when additional verification or skepticism is required.
Explainability and reasoning transparency “More visibility into the inner thought process.” (R1); “Clear reasoning steps and transparency about uncertainty.” (R19); “More transparent information about the thinking path.” (R44) Students can critically engage with AI outputs and treat AI as a cognitive support tool rather than an unquestioned authority.
Hallucination awareness and error signaling “Flagging hallucinatory content.” (R6); “Reduce hallucinations.” (R8); “Clear uncertainty indicators and fact-checking.” (R45) Students can identify potentially unreliable content and avoid incorporating incorrect information into academic work.
User feedback and human oversight “Letting you give them feedback, and use that to train in real-time.” (R38); “Allow users to flag hallucinated or incorrect information.” (R53) Supports human-in-the-loop interaction and increases accountability in educational AI systems.
Skepticism toward AI use “Features that prioritize perceived trustworthiness undermine nuance of AI deception.” (R11) Highlights the need for AI systems that respect skepticism rather than attempting to enforce trust or authority.
A dominant theme across participants’ open-ended responses concerns the need for transparency and verifiability in AI-generated outputs. Many participants emphasized that AI systems should consistently provide explicit source citations, clickable references, or other mechanisms that allow users to trace where information originates. Rather than treating AI outputs as authoritative answers, participants repeatedly framed trustworthy AI as a system that enables independent verification. As one participant explained, “Always returning a clickable link to a source so it can be easily validated” (R10). Similarly, another participant stated, “I think AI systems should always provide where it’s getting its data from so the user is able to check it on their own” (R42). These responses reflect a shared expectation that AI systems should support epistemic accountability by making their informational grounding visible, particularly in academic contexts where correctness and traceability are critical.
Closely related to source transparency is the expectation that AI systems should explicitly communicate uncertainty. Participants frequently requested confidence scores, reliability metrics, or clear indicators when the system is unsure or unable to answer. Several participants suggested that AI should not attempt to mask uncertainty through fluent or confident language. For example, one participant noted, “It should have a percentage next to every response. That percentage should tell how factual AI thinks the information that it presented is” (R23), while another stated, “I think I would like if it just shares when it’s not able to answer” (R2). Importantly, some participants cautioned that confident presentation can be misleading in educational settings. As R11 argued, “Features that prioritize perceived trustworthiness undermine the nuance of AI deception,” emphasizing that productive AI use requires sustained skepticism rather than blind trust. Together, these responses suggest that transparency about uncertainty is viewed as a prerequisite for responsible and effective AI use in learning contexts.
Explainability and reasoning transparency also emerged as a central design consideration. Participants expressed interest in greater visibility into how AI systems generate responses, including access to reasoning steps, intermediate logic, or the ability to intervene during response generation. One participant requested “more visibility into the inner thought process, more ability to do surgery on previous responses in a conversation” (R1), while another highlighted the importance of “clear reasoning steps, transparency about uncertainty, and real-time fact-checking” (R19). Several participants framed explainability not as a way to increase trust, but as a means of supporting critical engagement. For example, a participant described valuing systems that “show the steps of their thinking where I can intervene in the middle and reshape the direction.” These responses indicate that students prefer AI systems that function as interactive reasoning partners rather than opaque answer generators.
Participants also raised strong concerns about hallucinations and misleading outputs, underscoring the need for explicit error signaling and safeguards. Many respondents requested features that flag potentially hallucinated content, provide warnings, or incorporate built-in fact-checking mechanisms. As one participant succinctly stated, “Flagging hallucinatory content. Proper warnings (not in fine print) also help” (R6). Others emphasized reducing hallucinations altogether, with one participant stating, “Reduce hallucinations” (R8), and another calling for “clear uncertainty indicators, citations linked to verifiable sources, and explanations of reasoning steps” to increase trust (R45). These responses highlight participants’ awareness of AI failure modes and their desire for systems that proactively surface limitations rather than obscuring them.
Finally, a minority of participants expressed explicit skepticism toward the use of AI in education. Rather than attempting to impose trust through persuasive or opaque design choices, participants emphasized that effective educational use of AI requires acknowledging its limitations and potential for error. As one participant succinctly reflected, “I think features that prioritize perceived trustworthiness undermine the nuance of AI deception. It seems more pragmatic to me for educational perspectives to accept that AI is a tool that typically requires some skepticism to leverage effectively.” (R11)
6. Discussion
Across RQ1–RQ3, our findings collectively reveal that students’ engagement with AI in academic contexts is neither uniformly enthusiastic nor uniformly cautious. Instead, students demonstrate a nuanced, task-sensitive reasoning process that balances desired automation, actual usage, perceived benefits, and anticipated risks. In this section, we synthesize findings from all three research questions to discuss (1) how gaps between desired and actual automation reflect students’ conditional trust in AI systems (RQ1), (2) how motivations and concerns are selectively activated based on task characteristics (RQ2), and (3) how students’ articulated design expectations respond directly to these tensions (RQ3). Together, these insights highlight important implications for the design of educational AI systems that aim to support learning without undermining student agency.
6.1. Conditional Automation and the Limits of Adoption
Findings from RQ1 show a consistent gap between students’ desired levels of automation and their reported AI usage across most academic tasks. Importantly, this gap does not reflect general resistance to AI, but rather a pattern of conditional adoption. Students express openness to greater AI involvement while simultaneously withholding full reliance in practice.
This conditionality is strongly shaped by task characteristics. Tasks that are procedural, supportive, or easily verifiable (e.g., writing revision, summarization, citation formatting, and debugging code) tend to cluster in regions of both high desired automation and high usage. In contrast, tasks that are socially sensitive (e.g., drafting emails) or cognitively generative (e.g., brainstorming, quantitative reasoning) exhibit larger gaps, with desired automation exceeding actual usage. These patterns suggest that students do not evaluate AI solely based on its technical capability, but also based on contextual factors such as accountability, learning value, and error detectability.
From a broader perspective, RQ1 indicates that students are actively negotiating the boundary between assistance and delegation. Rather than seeking maximal automation, students appear to prefer retaining human control in tasks where outcomes carry social, intellectual, or evaluative significance. This finding challenges narratives that frame AI adoption as a linear progression toward increasing automation, and instead positions student–AI interaction as a calibrated and situational process.
6.2. Task-Specific Motivations and Risk Awareness
Results from RQ2 provide insight into why such conditional adoption emerges. Across tasks, efficiency-related motivations—particularly saving time and reducing cognitive load—dominate students’ reported reasons for using AI. However, these motivations are not applied uniformly. Instead, they interact with task-specific concerns that reflect students’ awareness of cognitive and epistemic risks.
For writing, revision, and learning support tasks, AI is primarily conceptualized as a pragmatic assistant. These tasks are characterized by outputs that are inspectable, editable, and low risk, which helps explain why concerns are minimal and why “None” frequently emerges as the dominant concern. In these contexts, students appear comfortable leveraging AI to streamline effort without feeling that core learning is compromised.
In contrast, ideation and technical problem-solving tasks activate distinct concern profiles. For brainstorming, concerns about reduced critical thinking dominate overwhelmingly, suggesting that students view ideation as central to intellectual ownership and learning. For technical and quantitative tasks, concerns about inaccurate information and shallow understanding feature prominently, reflecting the higher stakes associated with errors and the difficulty of verification. These findings demonstrate that students’ concerns are not abstract fears about AI, but grounded assessments of how AI assistance intersects with task goals and learning outcomes.
Taken together, RQ2 shows that students are not merely efficiency seekers. They are sensitive to the cognitive role of the task and actively evaluate whether AI use supports or threatens their learning objectives. This sensitivity provides important context for understanding the adoption gaps observed in RQ1.
6.3. Design Expectations as Responses to Identified Risks
Findings from RQ3 reveal that students’ expectations for AI system features directly respond to the tensions identified in RQ1 and RQ2. Rather than requesting more automation, participants consistently called for design features that support verification, reflection, and skepticism.
Transparency and verifiability emerged as the most dominant design expectation. Students repeatedly emphasized the need for source citations, clickable references, and traceable evidence, particularly in response to concerns about inaccurate information. These requests align closely with the elevated concern profiles observed in technical and writing-related tasks, where correctness is essential and errors may propagate easily.
Similarly, explicit communication of uncertainty—through confidence scores, reliability indicators, or admission of inability to answer— addresses students’ reluctance to over-rely on AI in cognitively demanding tasks. Rather than equating confidence with trust, students explicitly rejected persuasive or authoritative presentation, arguing instead for systems that make uncertainty visible. This stance reflects a mature understanding of AI limitations and positions skepticism as a productive component of educational AI use.
Explainability and opportunities for user intervention further reinforce this orientation. Students expressed interest in AI systems that expose reasoning steps and allow mid-course correction, not to eliminate effort, but to support active engagement and learning. Notably, even explicitly skeptical participants did not reject AI outright; instead, they emphasized the importance of designs that respect user agency and avoid creating an illusion of infallibility.
6.4. Implications for Educational AI Design
Synthesizing across all three research questions, our findings suggest that effective educational AI systems should prioritize calibrated assistance rather than maximal automation. Students’ desired future use of AI is not limited by lack of interest, but by concerns about trust, learning, and accountability.
Designs that foreground transparency, uncertainty communication, and explainability may help close the gap between desired and actual AI usage observed in RQ1, particularly for tasks where students currently hesitate. More broadly, our results indicate that supporting learning requires AI systems that scaffold thinking, invite verification, and preserve students’ role as primary decision-makers. Educational AI that emphasizes efficiency alone risks undermining precisely those cognitive processes students seek to protect.
7. Limitations
This study has several limitations that should be considered when interpreting the results.
First, the sample size of this study is relatively small and drawn from a single disciplinary population, namely computer science (CS) students. While CS students are typically early adopters of AI tools and are more familiar with interacting with LLM-based systems, their experiences and expectations may not generalize to students from other academic disciplines. Students in the humanities, social sciences, or professional fields may engage with AI differently, particularly in terms of task types, norms of academic integrity, and tolerance for automation. As a result, the patterns observed in desired automation, usage levels, and concerns may reflect discipline-specific practices rather than universal student perspectives. Future work should extend this investigation to a more diverse student population to examine how disciplinary context shapes student–AI interaction.
Second, unlike prior work (Shao et al., 2025), which incorporates expert annotation to assess the objective capabilities of AI systems on specific tasks, our study relies on self-reported measures of AI usage and perceived automation levels. This means that our analysis reflects students’ subjective experiences and judgments rather than externally validated performance metrics. While this approach is well suited for understanding students’ perceptions, trust calibration, and decision-making, it does not allow us to directly compare perceived AI capability with actual system performance. Consequently, mismatches between desired automation and usage should be interpreted as reflecting students’ beliefs and caution, rather than definitive assessments of AI effectiveness. Future research could combine user-reported data with expert evaluation or benchmark-based performance assessments to more fully characterize the relationship between perceived and actual AI capabilities in educational contexts.
Despite these limitations, our study provides valuable insights into how students reason about AI use across academic tasks, highlighting the nuanced trade-offs they make between efficiency, learning, and risk.
8. Conclusion
In this paper, we examined how graduate CS students engage with AI systems across a range of academic tasks, focusing on desired automation levels, actual usage, underlying motivations, perceived concerns, and expectations for AI system design. Through a mixed-methods user study, we showed that students’ adoption of AI is highly task dependent, with consistent gaps between desired and actual automation that reflect cautious, evaluative decision-making rather than indiscriminate use. We further demonstrated that efficiency-driven motivations coexist with task-specific concerns, particularly around inaccurate information and reduced critical thinking. Finally, we identified key design expectations for educational AI systems, including transparency, verifiability, uncertainty communication, and support for user skepticism. Together, our findings provide empirical grounding and design-relevant insights for building educational AI systems that align with students’ learning goals and preserve human agency in student–AI collaboration.