### v3

# Auditing Student–AI Collaboration: A Case Study of Online Graduate CS Students

Abstract.
As generative AI becomes embedded in higher education, it increasingly shapes how students complete academic tasks. While these systems offer efficiency and support, concerns persist regarding over-automation, diminished student agency, and the potential for unreliable or hallucinated outputs. This study conducts a mixed-methods audit of student–AI collaboration preferences by examining the alignment between current AI capabilities and students’ desired levels of automation in academic work. Using two sequential and complementary surveys, we capture students’ perceived benefits, risks, and preferred boundaries when using AI. The first survey employs an existing auditing framework to assess preferences for and actual usage of AI across 12 tasks in educational contexts, alongside primary concerns and reasons for use. The second survey, informed by the first, explores how AI systems could be designed to address these concerns through open-ended questions. This study aims to identify gaps between existing AI affordances and students’ normative expectations of collaboration, informing the development of more effective and trustworthy AI systems for education.
AI in education, human–AI collaboration, student agency, automation preferences, generative AI
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: Proceedings of the Bilaign Workshop at CHI 2026; April 2026; Barcelona, Spainbooktitle: Proceedings of the BiAlign Workshop at CHI 2026isbn: 978-1-4503-XXXX-X/26/06ccs: Human-centered computing Empirical studies in HCI
1. Introduction
Generative artificial intelligence (GenAI) tools such as ChatGPT, Claude, and Gemini are rapidly becoming embedded in higher education, reshaping how students complete academic tasks (Kasneci et al., 2023; Chan and Hu, 2025; Sukumar et al., 2025). These systems promise substantial gains in efficiency, personalization, and support, yet they also introduce significant risks related to over-automation, reduced student agency, and vulnerability to unreliable or hallucinated outputs (Shneiderman, 2020; Ji et al., 2023; Athaluri et al., 2023). As students increasingly delegate cognitive and procedural work to AI, it becomes essential to understand not only how they use these tools, but also how they perceive the trade-offs between automation and autonomy, and what design features would make AI systems more trustworthy in educational contexts (Khosravi et al., 2022; Amershi et al., 2019).
Recent research highlights the dual-edged nature of GenAI in education. Systematic reviews indicate that these tools can enhance writing quality, support personalized learning, and increase productivity, while simultaneously raising concerns about academic integrity, bias, and equitable access (Sukumar et al., 2025; Li et al., 2024; Camacho-Zúñiga et al., 2025). Empirical studies further suggest that the impact of AI depends heavily on how it is integrated pedagogically (Lee and Moore, 2024; Maizatul Akmar et al., 2025). However, a critical gap remains in understanding how students themselves navigate collaboration with AI across diverse tasks in educational setting.
This study addresses that gap by conducting a mixed-methods audit of student–AI collaboration among graduate students in the Georgia Tech Online Master of Science in Computer Science (OMSCS) program. Grounded in a Human-Centered AI (HCAI) perspective (Shneiderman, 2020, 2022) and informed by the Human Agency Scale (HAS) (Shao et al., 2025), we investigate two core dimensions of this collaboration. First, we measure students’ preferred levels of automation across twelve common tasks in educational contexts and compare these preferences to their actual AI usage (Shao et al., 2025, 2024b). This allows us to map tasks into four alignment zones: Green Light (high desire, high use), R&D Opportunity (high desire, low use), Low Priority (low desire, low use), and Red Light (low desire, high use). Second, through qualitative analysis of open-ended responses, we identify the system-level features that students believe would enhance the trustworthiness of AI in education (Amershi et al., 2019; Khosravi et al., 2022, 2024).
In this paper, we investigate the following questions:
  • RQ1: What levels of automation do students desire across different academic tasks, and how do these preferences align with their actual AI usage?
  • RQ2: How do students’ reasons for using AI and their associated concerns vary across task types, and what does this reveal about the perceived risks and benefits of student–AI collaboration?
  • RQ3: What system features and design principles do students identify as most important for addressing their concerns?
2. Related Work
2.1. Generative AI in Higher Education
A growing body of literature documents the transformative but contested role of generative AI in academic settings. Systematic reviews and meta-analyses highlight that tools such as ChatGPT can improve writing fluency, reduce cognitive load for routine tasks, and support personalized learning experiences when used appropriately (Kasneci et al., 2023). These benefits are most apparent for surface-level tasks like grammar correction, summarization, and initial drafting, where AI functions as a productivity aid rather than a substitute for learning.
At the same time, substantial concerns have been raised about over-reliance on AI, hallucinated or biased outputs, and the potential erosion of critical thinking and disciplinary epistemic skills (Bastani et al., 2024; Selwyn, 2019). Empirical work suggests that when students delegate substantive intellectual work—such as complex problem solving or argumentation—to AI systems without appropriate scaffolding, learning outcomes can suffer (Bastani et al., 2024).
Instructional context and pedagogical design play a decisive role in mediating these effects. For example, research comparing AI-generated feedback with human or peer feedback finds that AI is effective for addressing surface-level concerns aligned with rubrics, but that peer or instructor feedback remains essential for fostering deeper conceptual understanding and disciplinary reasoning (Cho and Schunn, 2007; Liu et al., 2024). Such work converges on the conclusion that generative AI is most beneficial when integrated as a complement to human pedagogical interaction, rather than as a replacement.
2.2. Human-Centered AI in Educational Contexts
Human-Centered AI (HCAI) has emerged as a foundational paradigm for responding to the risks posed by opaque, overly autonomous AI systems. Rooted in human–computer interaction and AI ethics, HCAI advocates for systems that are reliable, safe, and trustworthy while preserving meaningful human control and agency (Shneiderman, 2020). Rather than optimizing for automation alone, HCAI emphasizes augmentation that supports human goals, values, and accountability.
In educational contexts, HCAI principles foreground the importance of preserving both student and instructor agency, enhancing transparency of system behavior, and embedding safeguards against misuse, bias, and automation bias (Amershi et al., 2019; Holmes et al., 2022). Design recommendations from this literature include making system limitations visible, offering explanations and confidence cues, and enabling users to intervene, override, or verify AI outputs. These features are particularly critical in learning environments where inappropriate trust or over-dependence on AI can undermine motivation, metacognition, and skill development.
2.3. Human Agency and Automation Preferences
As AI systems become active collaborators in human workflows, researchers have sought to quantify how users perceive and negotiate their agency under automation. A recent contribution to this literature is the Human Agency Scale (HAS) introduced by Shao et al., which provides a five-level scale (H1–H5) to quantify the degree of human involvement preferred in task performance when working with AI agents (Shao et al., 2025). HAS moves beyond binary automate-or-not frameworks by offering a shared language for examining desired human control versus automation or augmentation across diverse tasks and occupations.
Our study adapts the HAS framework to education domain, applying it at a task-specific level to compare desired human agency and actual AI use. This approach enables a fine-grained audit of student–AI collaboration and provides empirical grounding for human-centered design recommendations.
3. Methodology
Table 1. 12 tasks in educational contexts used in the study.
ID Task Description
T1 Summarize a research article.
T2 Formatting Citation and Bibliography.
T3 Brainstorm essay ideas or outline structure
T4 Revise Writing for grammar and style.
T5 Debug code snippet or explain errors.
T6 Create personalized study plan.
T7 Generate flashcards or practice quizzes
T8 Recommend Learning Resources such as video or textbook
T9 Draft professional emails to TA or Professor
T10 Help explain or solve step by step quantitative problem
T11 Summarize class note or discussion
T12 Give feedback on draft writing
3.1. 12 Tasks in Educational Context
To systematically evaluate student preferences for AI collaboration, we selected twelve tasks in educational setting. Each task was chosen based on its relevance to both conceptual and practical skill development. For instance, tasks such as debugging code (T5) and solving quantitative problems step-by-step (T10) represent core technical competencies, while activities like summarizing research (T1) and providing feedback on writing (T12) emphasize critical engagement and communication skills (Shao et al., 2024a). By including tasks that range from routine (e.g., formatting citations) to complex and open-ended (e.g., brainstorming essay structures), we aimed to capture a spectrum of automation preferences and perceived risks.
Table 1 presents the complete list of tasks.
3.2. Desired vs. Actual Automation Likert Questions
Likert Question for Collecting Automation Desire Question. For each academic task, please rate the level of automation that an AI system should ideally provide to support learning, from your (a student’s) perspective. Scale. 1: AI should not get involved in this task at all 2: AI may give minor suggestions 3: AI and human should share responsibility 4: AI should handle most of the task with human oversight 5: AI should perform this task autonomously
Figure 1. Desired level of AI automation.
Likert Question for Collecting Observed AI Usage Question. For each academic task listed below, please indicate how capable current AI systems are based on your own experience, reflected by how you actually use AI for that task. Scale. 1: I do not use AI for this task at all 2: I use AI to provide minor or partial help 3: I use AI for this task moderately 4: I use AI to handle most of the task with oversight 5: I rely entirely on AI to automate this task
Figure 2. Actual AI usage Level.
Following prior work by Shao et al. (Shao et al., 2025) on characterizing levels of AI automation, we operationalize students’ automation preferences using a five-point Likert-scale that reflects a continuum of responsibility allocation between humans and AI (Figure 1). Rather than treating automation as a binary choice, this scale explicitly distinguishes between AI serving a purely assistive role, sharing responsibility with the human learner, or assuming primary or full control over the task (Shao et al., 2024b). This design allows us to capture nuanced automation boundaries across academic tasks and to examine where students draw limits on AI involvement in educational contexts (Venkatesh et al., 2003).
In contrast to prior work by Shao et al. (Shao et al., 2025), which characterizes AI capability using expert annotations, our study does not rely on external expert judgments of system performance. Instead, we capture students’ perceived AI capability through their self-reported actual usage and reliance on AI systems across academic tasks (Figure 2(Dwivedi et al., 2023; Sullivan et al., 2023).
Specifically, we ask participants to indicate how they currently use AI for each task, ranging from not using AI at all to relying entirely on AI to automate the task. This formulation operationalizes perceived capability in terms of habitual use and practical reliance, reflecting students’ lived experiences with contemporary AI tools rather than an objective assessment of model correctness or task success (Jian et al., 2000).
3.3. Four-Zone Classification of AI Automation
To analyze students’ perceptions of appropriate AI involvement in academic tasks, we adopted a four-zone classification of automation adapted from Shao et al.’s framework (Shao et al., 2025) for integrating human worker and AI expert perspectives on automation. This framework partitions the automation landscape into four zones based on the alignment or misalignment between human preferences and perceived AI capability.
The four zones are defined as follows:
Automation “Green Light” Zone. Tasks in the Green Light zone are those for which both humans and AI systems are perceived as well-suited for automation. In this zone, there is strong alignment between participants’ willingness to delegate a task to AI and their confidence in AI’s ability to perform the task effectively. These tasks represent contexts in which AI adoption is broadly acceptable and likely to provide immediate benefits with minimal resistance (Shao et al., 2025).
Automation “Red Light” Zone. The Red Light zone includes tasks that participants believe should not be automated, even if AI systems are technically capable of performing them. This zone reflects strong human resistance to AI involvement, often due to concerns related to trust, accountability, ethical implications, or academic integrity (Cotton et al., 2023; Perkins, 2023). Tasks in this zone signal boundaries where AI use is perceived as inappropriate despite potential performance gains.
R&D Opportunity Zone. Tasks in the R&D Opportunity zone are those for which participants express openness to AI assistance but lack confidence in current AI systems’ capabilities. This zone highlights gaps between user expectations and existing technological performance and points to opportunities for further AI development, evaluation, and design refinement (Shao et al., 2025). Improving transparency, reliability, or controllability of AI systems may enable tasks in this zone to transition toward the Green Light zone (Kasneci et al., 2023).
Low Priority Zone. The Low Priority zone consists of tasks that participants neither desire to automate nor believe AI systems are well-suited to perform. These tasks are typically viewed as either unimportant for automation or inherently human-centered, resulting in low perceived value of AI involvement. From a design perspective, this zone suggests limited benefit in prioritizing AI development for such tasks (Shao et al., 2025).
3.4. Primary Concerns and Reasons for AI Usage across Academic Tasks
Category Survey Item Description
Primary Concerns
Inaccurate or misleading information Concern that AI outputs may contain factual errors or hallucinations that affect task correctness
Risk of academic misconduct Concern about plagiarism, policy violations, or inappropriate AI use in graded work
Reduced critical thinking Concern that relying on AI may weaken students’ independent reasoning or learning
Reasons for AI Usage
Saving time Using AI to complete tasks more efficiently and reduce time spent on repetitive work
Reducing cognitive load Using AI to manage mentally demanding or complex tasks
Improving output quality Using AI to enhance clarity, accuracy, or polish of academic outputs
Table 2. Selection of primary concerns and reasons for AI usage included in the survey.
To understand what shapes students’ adoption of AI across different academic tasks, we asked participants to report both their primary concerns when using AI and their reasons for relying on AI assistance (Dwivedi et al., 2023; Sullivan et al., 2023). Concerns included risks such as inaccurate or misleading outputs, hallucinations, academic misconduct, and reduced critical thinking (Cotton et al., 2023; Halaweh, 2023; Imran and Almusharraf, 2023), while reasons for use captured perceived benefits such as saving time, reducing cognitive load while managing conceptually complex task and improving accuracy or polishing academic outputs (Kasneci et al., 2023), as shown in Table 2.
3.5. Design Expectations for Addressing AI Concerns
In addition to structured Likert-scale items, we included open-ended question to elicit students’ perspectives on how AI systems should be designed to address their greatest concern in educational contexts (Braun and Clarke, 2006). The use of open-ended questions allows participants to express design expectations in their own terms, capturing aspects of AI system behavior and interface design that may not be fully anticipated by closed-form survey items (Braun and Clarke, 2006). We treat these responses as qualitative inputs for identifying recurring design considerations and informing subsequent analysis of user-centered approaches to AI alignment in education (Swiecki et al., 2022).
4. User Study Method
4.1. Participants
Participants were graduate students enrolled in the Online Master of Science in Computer Science (OMSCS) program at the Georgia Institute of Technology. Participation in the survey was voluntary. To ensure the relevance of responses to AI-assisted coursework, we filtered out participants who reported rarely using or not using AI tools in their academic work. In total, $N=57$ students completed the survey. After filtering, $N=44$ participants who reported at least occasional use of AI systems for coursework were retained for analysis. The final sample therefore consists of students with meaningful experience interacting with AI in educational contexts, allowing us to focus our analysis on substantive patterns of AI use.
4.2. Survey Based Interviews
In Phase 1, participants completed a questionnaire assessing their use of AI tools and their perceptions of appropriate levels of AI automation across a range of tasks in educational contexts. Participants also reported their primary reasons and concerns regarding AI use for each task. Using the responses from Phase 1, we identified the most frequently reported concern across tasks through quantitative aggregation of participants’ ratings. In Phase 2, we conducted a follow-up survey focused specifically on this most prevalent concern. Participants were asked to describe the AI system features, design principles, or safeguards they believed would effectively address this concern in educational contexts. This second phase enabled us to move from identifying perceived risks to systematically collecting participant-informed design expectations.
4.3. Ethics
The study followed established ethical guidelines for research involving human participants. Prior to participation, all participants completed an informed consent form. Participation was voluntary, and no personally identifiable information was collected. All responses were anonymized prior to analysis and used solely for research purposes.
The study posed minimal risk to participants. All questions focused on participants’ experiences and perceptions of AI use in academic contexts, and no deception was involved. Data were collected and stored in a manner that protected participant privacy and ensured confidentiality.
4.4. Data Analysis
To compare students’ desired levels of AI involvement with their reported use of AI tools, we operationalized two aggregate measures: aggregate desire level and aggregate usage level. Both measures were derived from participants’ Likert-scale responses and computed at the task level.
Aggregate Desire Level.
For each academic task $t$, participants indicated the level of AI automation they believed was appropriate using an ordered Likert-scale corresponding to increasing levels of automation (from minimal AI involvement to fully autonomous AI execution). Responses were numerically encoded such that higher values indicate a greater desired level of AI automation. The aggregate desire level for task $t$ was calculated as the mean of these encoded responses across all participants in the analytic sample ($N=44$):
$D_{t}=\frac{1}{N}\sum_{i=1}^{N}d_{i,t},$
where $d_{i,t}$ denotes participant $i$’s desired automation rating for task $t$.
Aggregate Usage Level.
Actual AI usage was operationalized using participants’ self-reported frequency of AI tool use. Responses were encoded on an ordered numerical scale, with higher values indicating more frequent AI use. The aggregate usage level was computed as the mean usage score across all participants and tasks:
$U_{t}=\frac{1}{N}\sum_{i=1}^{N}u_{i,t}.$
5. User Study Results
Figure 3. Comparison of students’ desired AI automation levels and self-reported actual AI usage across twelve tasks (T1–T12).
Figure 4. Four-zone automation alignment map plotting desired AI automation against actual AI usage for twelve tasks.
5.1. RQ1: Desired Automation Level vs. Actual AI Usage Level
Figure 3 illustrates participants’ aggregate desired automation levels and reported AI usage levels across twelve academic tasks (T1–T12). Across most tasks, desired automation levels exceed reported usage, indicating a consistent gap between participants’ preferences for AI involvement and their current patterns of AI use. However, the magnitude of this gap varies substantially by task type, suggesting that task characteristics play an important role in shaping how students engage with AI tools in practice.
5.1.1. Information Compression and Retrieval Tasks
Summarization and information organization (T1, T11) show relatively high desired automation and moderately high usage. These largely extractive, low-reasoning tasks are perceived as supportive rather than generative. Because errors are easier to detect and consequences are limited, participants are comfortable using AI, resulting in a relatively small desire–usage gap.
5.1.2. Writing Support and Surface-Level Editing Tasks
Grammar revision (T4), citation formatting (T2), and draft feedback (T12) exhibit some of the highest desired automation levels. As procedural and rule-based activities, they are well suited to AI assistance. However, actual usage remains lower than desired, suggesting caution related to correctness, formatting standards, and academic norms.
5.1.3. Ideation and Study Support Tasks
Brainstorming (T3), study planning (T6), flashcard generation (T7), and resource recommendation (T8) display some of the largest desire–usage gaps. Although participants express strong interest in AI support, usage is comparatively low. These tasks require contextualization and personalization, areas where AI is perceived as inconsistently reliable, producing openness tempered by reservations.
5.1.4. Communication and Socially Sensitive Tasks
Professional email drafting (T9) shows moderate desired automation but relatively low usage. Here, the gap appears driven less by perceived capability and more by interpersonal concerns. Participants may hesitate to delegate tasks involving tone, professionalism, and accountability.
5.1.5. Analytical and High-Stakes Reasoning Tasks
Analytical tasks vary. For code debugging (T5), usage closely aligns with desired automation, reflecting confidence in outputs that can be tested. In contrast, quantitative problem solving (T10) shows a larger gap. Despite moderate interest in AI support, participants report lower usage, likely due to the precision required and the difficulty of detecting reasoning errors.
5.1.6. Four Automation Zones
Figure 4 shows that the twelve academic tasks are distributed unevenly across the four automation zones, revealing systematic patterns in how students’ preferences for AI automation align with their reported usage.
Several tasks cluster in the region characterized by both relatively high desired automation and high actual AI usage. These include formatting citations and bibliographies, revising writing for grammar and style, summarizing class notes or discussions, summarizing research articles, and debugging code. The concentration of tasks in this region suggests that students are already actively using AI for tasks that are procedural, supportive, or whose outputs can be readily verified.
A second group of tasks falls into a region where desired automation is high but actual usage remains comparatively lower. This group includes creating personalized study plans, generating flashcards or practice quizzes, recommending learning resources, drafting professional emails, and solving step-by-step quantitative problems. The placement of these tasks indicates that while students express interest in greater AI involvement, their current usage lags behind their preferences, highlighting areas where existing AI tools may not yet fully meet students’ expectations or where contextual concerns limit adoption.
5.2. RQ2: Primary Reasons and Concerns of using AI across tasks
Figure 5. Distribution of students’ reported reasons for using AI across tasks.
Figure 6. Distribution of students’ reported concerns regarding AI use across tasks.
Figures 6 illustrate the distributions of students’ reported reasons for using AI, and their associated concerns across twelve academic tasks was illustrated in Figures 6.
5.2.1. Writing and Revision Tasks
Writing- and revision-oriented tasks (T1, T4, T9, T11, T12) are strongly dominated by efficiency-related motivations. Across this category, saving time accounts for the largest share of responses, often exceeding half. Students primarily view AI as pragmatic assistance that reduces routine effort rather than reshapes intellectual substance. Improving output quality is a substantial secondary reason, particularly for revision tasks (T4, T12), where AI enhances clarity and surface-level correctness.
These patterns suggest that students see AI as complementing—not replacing—their writing. Because outputs are inspectable and editable, barriers to adoption are low, and students retain control over content and intent.
Concerns are comparatively limited. For several tasks (e.g., T9, T11), None is the most common response. When concerns arise, they focus mainly on inaccurate information, outweighing worries about plagiarism or reduced critical thinking. Overall, trust in AI writing support appears conditional on students’ ability to verify outputs.
5.2.2. Ideation and Planning Tasks
Ideation and planning tasks (T3, T6) show more differentiated patterns. For brainstorming (T3), motivations are distributed across improving output quality and reducing cognitive load, reflecting interest in idea generation and overcoming initial barriers. However, reduced critical thinking overwhelmingly dominates concerns, indicating awareness of the cognitive risks of delegating creative reasoning. Ideation thus represents a sensitive boundary between assistance and substitution.
In contrast, study planning (T6) presents a simpler profile: saving time is the dominant reason, and None is the most common concern. Students appear to view planning as organizational rather than cognitively generative, making AI use less threatening to learning.
5.2.3. Learning Support Tasks
Learning support tasks (T7, T8) exhibit highly concentrated reason distributions, with saving time clearly dominant and improving output quality secondary. Students primarily view AI as accelerating access to study materials.
Concerns are minimal: None is the largest category, and accuracy- or cognition-related worries are limited. Because outputs such as flashcards or resource recommendations can be verified or selectively used, AI is perceived as a low-stakes, supplementary aid.
5.2.4. Technical and Quantitative Problem-Solving Tasks
Technical tasks (T5, T10) display the most balanced and complex profiles. Reducing cognitive load is the leading motivation, with improving output quality also prominent, reflecting the complexity and precision required.
At the same time, concerns remain salient. Inaccurate information and reduced critical thinking together account for a substantial share, exceeding other categories—particularly inaccurate information, which is the most frequently reported concern across tasks. Students value AI for scaffolding complex reasoning but remain cautious about errors and superficial understanding. Compared to writing or general learning tasks, mistakes in technical contexts are harder to detect and potentially more consequential, leading to a more evaluative and risk-sensitive stance toward AI use.
Table 3. Students’ expected features and design principles for educational AI systems.
Reason Example quotes Potential benefits
Transparency and verifiability of AI outputs “Always returning a clickable link to a source so it can be easily validated.” (R10); “Giving sources for all the information they give.” (R22); “AI systems should always provide where it’s getting its data from.” (R42) Students can independently verify AI-generated information and assess its reliability, reducing blind trust and inappropriate academic use.
Explicit communication of uncertainty “It should have a percentage next to every response.” (R23); “A certainty factor would be an interesting addition.” (R4); “More transparency about uncertainty or when guessing is used.” (R34) Students can better calibrate trust in AI outputs and decide when additional verification or skepticism is required.
Explainability and reasoning transparency “More visibility into the inner thought process.” (R1); “Clear reasoning steps and transparency about uncertainty.” (R19); “More transparent information about the thinking path.” (R44) Students can critically engage with AI outputs and treat AI as a cognitive support tool rather than an unquestioned authority.
Hallucination awareness and error signaling “Flagging hallucinatory content.” (R6); “Reduce hallucinations.” (R8); “Clear uncertainty indicators and fact-checking.” (R45) Students can identify potentially unreliable content and avoid incorporating incorrect information into academic work.
User feedback and human oversight “Letting you give them feedback, and use that to train in real-time.” (R38); “Allow users to flag hallucinated or incorrect information.” (R53) Supports human-in-the-loop interaction and increases accountability in educational AI systems.
Skepticism toward AI use “Features that prioritize perceived trustworthiness undermine nuance of AI deception.” (R11) Highlights the need for AI systems that respect skepticism rather than attempting to enforce trust or authority.
5.3. RQ3: Expected System Features and Design Principles of AI System
After participants reported their primary reasons and concerns regarding AI use across academic tasks, we administered a follow-up survey to collect open-ended responses about the design features they expected AI systems to provide to address the most prevalent concerns across tasks: inaccurate information. This resulted in 53 valid responses ($N=53$) out of the 57 participants, in which respondents described the features and design principles they expect AI systems to provide to address concerns about inaccurate information and hallucinations. We refer to individual respondents as R1–R53 throughout this section.
As Table 3 shows, a dominant theme across participants’ open-ended responses concerns the need for transparency and verifiability in AI-generated outputs. Many participants emphasized that AI systems should consistently provide explicit source citations, clickable references, or other mechanisms that allow users to trace where information originates. Rather than treating AI outputs as authoritative answers, participants repeatedly framed trustworthy AI as a system that enables independent verification. As one participant explained, “Always returning a clickable link to a source so it can be easily validated” (R10). Similarly, another participant stated, “I think AI systems should always provide where it’s getting its data from so the user is able to check it on their own” (R42).
Closely related to source transparency is the expectation that AI systems should explicitly communicate uncertainty. Participants frequently requested confidence scores, reliability metrics, or clear indicators when the system is unsure or unable to answer. Several participants suggested that AI should not attempt to mask uncertainty through fluent or confident language. For example, one participant noted, “It should have a percentage next to every response. That percentage should tell how factual AI thinks the information that it presented is” (R23), while another stated, “I think I would like if it just shares when it’s not able to answer” (R2). Importantly, some participants cautioned that confident presentation can be misleading in educational settings. As R11 argued, “Features that prioritize perceived trustworthiness undermine the nuance of AI deception,” emphasizing that productive AI use requires sustained skepticism rather than blind trust.
Explainability and reasoning transparency also emerged as a central design consideration. Participants expressed interest in greater visibility into how AI systems generate responses, including access to reasoning steps, intermediate logic, or the ability to intervene during response generation. One participant requested “more visibility into the inner thought process, more ability to do surgery on previous responses in a conversation” (R1), while another highlighted the importance of “clear reasoning steps, transparency about uncertainty, and real-time fact-checking” (R19). Several participants framed explainability not as a way to increase trust, but as a means of supporting critical engagement. For example, a participant described valuing systems that “show the steps of their thinking where I can intervene in the middle and reshape the direction.”
Participants also raised strong concerns about hallucinations and misleading outputs, underscoring the need for explicit error signaling and safeguards. Many respondents requested features that flag potentially hallucinated content, provide warnings, or incorporate built-in fact-checking mechanisms. As one participant succinctly stated, “Flagging hallucinatory content. Proper warnings (not in fine print) also help” (R6). Others emphasized reducing hallucinations altogether, with one participant stating, “Reduce hallucinations” (R8), and another calling for “clear uncertainty indicators, citations linked to verifiable sources, and explanations of reasoning steps” to increase trust (R45).
6. Discussion
6.1. Conditional Automation and the Limits of Adoption
Findings from RQ1 show a consistent gap between students’ desired levels of automation and their reported AI usage across most academic tasks. Importantly, this gap does not reflect general resistance to AI, but rather a pattern of conditional adoption. Students express openness to greater AI involvement while simultaneously withholding full reliance in practice.
This conditionality is strongly shaped by task characteristics. Tasks that are procedural, supportive, or easily verifiable (e.g., writing revision, summarization, citation formatting, and debugging code) tend to cluster in regions of both high desired automation and high usage. In contrast, tasks that are socially sensitive (e.g., drafting emails) or cognitively generative (e.g., brainstorming, quantitative reasoning) exhibit larger gaps, with desired automation exceeding actual usage. These patterns suggest that students do not evaluate AI solely based on its technical capability, but also based on contextual factors such as accountability, learning value, and error detectability.
6.2. Task-Specific Motivations and Risk Awareness
Results from RQ2 provide insight into why such conditional adoption emerges. Across tasks, efficiency-related motivations—particularly saving time and reducing cognitive load—dominate students’ reported reasons for using AI. However, these motivations are not applied uniformly. Instead, they interact with task-specific concerns that reflect students’ awareness of cognitive and epistemic risks.
For writing, revision, and learning support tasks, AI is primarily conceptualized as a pragmatic assistant. These tasks are characterized by outputs that are inspectable, editable, and low risk, which helps explain why concerns are minimal and why “None” frequently emerges as the dominant concern. In these contexts, students appear comfortable leveraging AI to streamline effort without feeling that core learning is compromised.
In contrast, ideation and technical problem-solving tasks activate distinct concern profiles. For brainstorming, concerns about reduced critical thinking dominate overwhelmingly, suggesting that students view ideation as central to intellectual ownership and learning. For technical and quantitative tasks, concerns about inaccurate information and shallow understanding feature prominently, reflecting the higher stakes associated with errors and the difficulty of verification. These findings demonstrate that students’ concerns are not abstract fears about AI, but grounded assessments of how AI assistance intersects with task goals and learning outcomes.
6.3. Design Expectations as Responses to Identified Risks
Findings from RQ3 reveal that students’ expectations for AI system features directly respond to the tensions identified in RQ1 and RQ2. Rather than requesting more automation, participants consistently called for design features that support verification, reflection, and skepticism.
Transparency and verifiability emerged as the most dominant design expectation. Students repeatedly emphasized the need for source citations, clickable references, and traceable evidence, particularly in response to concerns about inaccurate information. These requests align closely with the elevated concern profiles observed in technical and writing-related tasks, where correctness is essential and errors may propagate easily.
Similarly, explicit communication of uncertainty—through confidence scores, reliability indicators, or admission of inability to answer— addresses students’ reluctance to over-rely on AI in cognitively demanding tasks. Rather than equating confidence with trust, students explicitly rejected persuasive or authoritative presentation, arguing instead for systems that make uncertainty visible. This stance reflects a mature understanding of AI limitations and positions skepticism as a productive component of educational AI use.
7. Conclusion
In this paper, we examined how graduate CS students engage with AI systems across a range of academic tasks, focusing on desired automation levels, actual usage, underlying motivations, perceived concerns, and expectations for AI system design. Through a mixed-methods user study, we showed that students’ adoption of AI is highly task dependent, with consistent gaps between desired and actual automation that reflect cautious, evaluative decision-making rather than indiscriminate use. We further demonstrated that efficiency-driven motivations coexist with task-specific concerns, particularly around inaccurate information and reduced critical thinking. Finally, we identified key design expectations for educational AI systems, including transparency, verifiability, uncertainty communication, and support for user skepticism. Together, our findings provide empirical grounding and design-relevant insights for building educational AI systems that align with students’ learning goals and preserve human agency in student–AI collaboration.
Appendix A Full Participant List
Table 4. Participant-level demographics of OMSCS students who reported AI tool use in coursework (N=57).
ID Age Group OMSCS Specialization AI Usage Frequency
P1 30–40 Human–Computer Interaction Several times per week
P2 25–30 Machine Learning Daily or almost daily
P3 25–30 Human–Computer Interaction Often
P4 18–24 Artificial Intelligence Often
P5 30–40 Human–Computer Interaction Several times per week
P6 25–30 Human–Computer Interaction Rarely
P7 30–40 Artificial Intelligence Several times per week
P8 Above 40 Computing Systems Several times per week
P9 30–40 Artificial Intelligence Daily or almost daily
P10 18–24 Artificial Intelligence Several times per week
P11 18–24 Not specified Daily or almost daily
P12 18–24 Artificial Intelligence Daily or almost daily
P13 18–24 Undecided Often
P14 30–40 Undecided Daily or almost daily
P15 30–40 Human–Computer Interaction Daily or almost daily
P16 30–40 Computing Systems Rarely
P17 18–24 Undecided Several times per week
P18 30–40 Computing Systems Often
P19 25–30 Artificial Intelligence Often
P20 30–40 Human–Computer Interaction Several times per week
P21 25–30 Computing Systems Several times per week
P22 18–24 Human–Computer Interaction Rarely
P23 18–24 Human–Computer Interaction Several times per week
P24 30–40 Artificial Intelligence Daily or almost daily
P25 18–24 Artificial Intelligence Often
P26 25–30 Human–Computer Interaction Rarely
P27 18–24 Human–Computer Interaction Rarely
P28 18–24 Human–Computer Interaction Rarely
P29 30–40 Machine Learning Daily or almost daily
P30 25–30 Human–Computer Interaction Rarely
P31 25–30 Human–Computer Interaction Daily or almost daily
P32 25–30 Computing Systems Daily or almost daily
P33 30–40 Artificial Intelligence Daily or almost daily
P34 18–24 Machine Learning Daily or almost daily
P35 25–30 Human–Computer Interaction Never
P36 30–40 Artificial Intelligence Never
P37 18–24 Artificial Intelligence / Machine Learning Several times per week
P38 Above 40 Artificial Intelligence Daily or almost daily
P39 18–24 Undecided Several times per week
P40 18–24 Human–Computer Interaction Rarely
P41 25–30 Artificial Intelligence Daily or almost daily
P42 18–24 Artificial Intelligence Daily or almost daily
P43 18–24 Not specified Daily or almost daily
P44 25–30 Human–Computer Interaction Rarely
P45 25–30 Human–Computer Interaction Several times per week
P46 18–24 Human–Computer Interaction Rarely
P47 18–24 Artificial Intelligence Often
P48 18–24 Artificial Intelligence Often
P49 30–40 Computing System Daily or almost daily
P50 25–30 Human–Computer Interaction Daily or almost daily
P51 25–30 Human–Computer Interaction Several times per week
P52 Above 40 Machine Learning Daily or almost daily
P53 Above 40 Human–Computer Interaction Rarely
P54 18–24 Artificial Intelligence Several times per week
P55 25–30 Human–Computer Interaction Several times per week
P56 25–30 Human–Computer Interaction Daily or almost daily
P57 30–40 Artificial Intelligence Daily or almost daily